Use the integral test to find whether the following series converge or diverge. Co (n2 + 1)2


Math Physics:

IMPORTANT! Step by step *typed* format preferred. If hand written, please write clearly and legibly. Please be sure answer is CORRECT! (No wrong answers please!) Thanks!
Answer #1

The series will converge if the following integral will converge,

1 (22+1)2

Evaluating the integral,

We will take,

t = x^2 + 1 \Rightarrow dt = 2xdx

The limits will change to 2 to infinity.

Substituting we get,

\int_2^\infty \frac{dt}{2t^2} = \frac{1}{2}\int_2^\infty \frac{dt}{t^2} = -\frac{1}{2} \left |\frac{1}{t} \right |^\infty _2 = -\frac{1}{2} \left(\frac{1}{\infty} - \frac{1}{2} \right ) = \frac{1}{4}

Since the integral converges to a value, therefore the above series converges to a value.